OpenStudy (anonymous):

PLEASE HELP ME Given the function f(x) = 4x^2 - |x - 12|, find f(5). 93 83 117 107

OpenStudy (anonymous):

wolframalpha.com or cymath.com

OpenStudy (anonymous):

\(\huge\frak\color{cyan}{I~Can~Help!!}\)

OpenStudy (anonymous):

:) !!

OpenStudy (anonymous):

What's your choices?

OpenStudy (anonymous):

93 83 117 107

OpenStudy (anonymous):

Example: Graph the function ƒ(x) = 2x + 5. Solution: Choose a value for x. Let x = 1 Substitute 1 in for x. ƒ(x) = 2(1) + 5 Simplify to find the value of y. ƒ(x) = 2 + 5 ƒ(x) = 7 Check. 7 2(1) + 5 7 2 + 5 7 = 7 (True) An ordered pair is (1, 7).

OpenStudy (anonymous):

To graph two variable equations we can find ordered pairs and plot them in the coordinate plane. To find an ordered pair, (x, y) : Pick a value for x. Substitute that value into the equation. Simplify to find the value of ƒ(x). Don't forget to use the order of operations! Check by substituting both values into the equation. State the ordered pair.

OpenStudy (anonymous):

put your information into the calculator and it will give you your answer:)

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

ok im back i was trying to figure it out but i dont know

OpenStudy (anonymous):

@Xx_H20_Avenger_xX

OpenStudy (anonymous):

I can't give direct answers..so you dont have choices? \(\huge\color{cyan}\heartsuit\)

OpenStudy (anonymous):

do have choices I posted them like twice

OpenStudy (anonymous):

ok so.... i know the answer and but i can't give direct answers

OpenStudy (anonymous):

i didnt ask for a direct answer....

OpenStudy (anonymous):

\(\huge\frak\color{#ff1ff1}{O...........K}\)

OpenStudy (anonymous):

someones mad.....

OpenStudy (anonymous):

\[\huge\color{cyan}\heartsuit\]